# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ Appendix C Congleton Formalism for Hypertriton
$\begin{pmatrix}\bf q_{1}\\ \mathbf{q}_{2}\\ \mathbf{P}\end{pmatrix}=J^{-1}\times\begin{pmatrix}\mathbf{p}_{1}\\ \mathbf{p}_{2}\\ \mathbf{p}_{3}\end{pmatrix}$ (84)
and
$\mathbf{p}_{1}$ $=\frac{\kappa(2+\kappa)\mathbf{P}}{2+\kappa^{2}}+\frac{(2+\kappa)\mathbf{q}_{2}}{2+\kappa^{2}}\,$ (85)
$\mathbf{p}_{2}$ $=\frac{(2+\kappa)\mathbf{P}}{2+\kappa^{2}}+\mathbf{q}_{1}-\frac{\kappa(2+\kappa)\mathbf{q}_{2}}{2(2+\kappa^{2})},$
$\mathbf{p}_{3}$ $=\frac{(2+\kappa)\mathbf{P}}{2+\kappa^{2}}-\mathbf{q}_{1}-\frac{\kappa(2+\kappa)\mathbf{q}_{2}}{2(2+\kappa^{2})},$
@@
satisfying the condition, $\mathbf{p}_{1}\cdot\mathbf{x}_{1}+\mathbf{p}_{2}\cdot\mathbf{x}_{2}+\mathbf{p}_{3}\cdot\mathbf{x}_{3}=\bm{\xi}_{1}\cdot\mathbf{q}_{1}+\bm{\xi}_{2}\cdot\mathbf{q}_{2}+\mathbf{R}\cdot\mathbf{P}$. Finally defining $\mathbf{R}=\frac{\bm{\zeta}_{1}+\bm{\zeta}_{2}}{2}$ and $\mathbf{R}^{\prime}=\frac{\bm{\zeta}_{2}-\bm{\zeta}_{1}}{2}$ as well as $\mathrm{d}^{3}x_{1}\,\mathrm{d}^{3}x_{2}\,\mathrm{d}^{3}x_{3}\,\mathrm{d}^{3}x_{1}^{\prime}\,\mathrm{d}^{3}x_{2}^{\prime}\,\mathrm{d}^{3}x_{3}^{\prime}=\frac{|J^{-1}|^{6}}{8}\,\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\,\mathrm{d}^{3}\xi_{1}^{\prime}\,\mathrm{d}^{3}\xi_{2}^{\prime}\,\mathrm{d}^{3}\zeta_{1}\,\mathrm{d}^{3}\zeta_{2}$ with $\mathrm{d}^{3}R\,\mathrm{d}^{3}R^{\prime}=\frac{1}{8}\,\mathrm{d}^{3}\zeta_{1}\,\mathrm{d}^{3}\zeta_{2}$. We obtain the yield in terms of the Jacobi coordinates and the Hypertriton wave function as
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{12}}\left(\frac{|J^{-1}|}{2}\right)^{3}\,\int\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\,\mathrm{d}^{3}\xi_{1}^{\prime}\,\mathrm{d}^{3}\xi_{2}^{\prime}\,\mathrm{d}^{3}\zeta_{1}\,\mathrm{d}^{3}\zeta_{2}\,\,\mathrm{d}^{3}q_{1}\,\mathrm{d}^{3}q_{2}\,\mathrm{d}^{3}P\,\,e^{i\bm{\zeta}_{1}\cdot(\mathbf{P}-\mathbf{P}_{{{}^{3}_{\Lambda}{\rm H}}})}e^{-\frac{\left(2+\kappa^{2}\right)\zeta_{2}^{2}}{8\sigma^{2}(2+\kappa)^{2}}}\\ \times\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\,\,e^{i(\mathbf{q}_{1}\cdot\bm{\xi}_{1}-\mathbf{q}_{1}\cdot\bm{\xi}_{1}^{\prime}+\mathbf{q}_{2}\mathbf{\cdot}\bm{\xi}_{2}-\mathbf{q}_{2}\cdot\bm{\xi}_{2}^{\prime})}\times(2\pi\sigma^{2})^{-9/2}e^{\chi^{2}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}$ (86)
@@
with a substitution of defined in terms of $\chi^{2}=-\frac{4\left((\xi_{1}+\xi_{1}^{\prime})^{2}+2(\xi_{2}+\xi_{2}^{\prime})^{2}\right)+4\kappa(\xi_{1}+\xi_{1}^{\prime})^{2}+\kappa^{2}\left(\xi_{1}^{2}+\bm{\xi}_{1}^{\prime}\cdot(2\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})+4\xi_{2}^{2}+4\bm{\xi}_{2}^{\prime}\cdot(2\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})\right)}{16\sigma^{2}(2+\kappa)^{2}}$. Performing pure $\delta$ integral $\int\mathrm{d}^{3}\zeta_{1}\,e^{i(\mathbf{P}-\mathbf{P}_{{{}^{3}_{\Lambda}{\rm H}}})\cdot\bm{\zeta}_{1}}=(2\pi)^{3}\delta(\mathbf{P}-\mathbf{P}_{{{}^{3}_{\Lambda}{\rm H}}})$ and $\int\mathrm{d}^{3}P$ and $\int\mathrm{d}^{3}\zeta_{2}$ in Eq. 86
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{9}(2\pi\sigma^{2})^{9/2}}\,\frac{8\sqrt{2}\pi^{3/2}{|J^{-1}|^{3}}\sigma^{3}(2+\kappa)^{3}}{\left(2+\kappa^{2}\right)^{3/2}}\,\int\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\,\mathrm{d}^{3}\xi_{1}^{\prime}\,\mathrm{d}^{3}\xi_{2}^{\prime}\,\,\mathrm{d}^{3}q_{1}\,\mathrm{d}^{3}q_{2}\,\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\\ \times e^{i(\mathbf{q}_{1}\cdot(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime})+\mathbf{q}_{2}\cdot(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}))}\times e^{\chi^{2}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}$ (87)
@@
Further simplifications are made by making the following coordinate transformations $\bm{\xi}_{1}=\mathbf{r}_{1}+\bm{\zeta}_{1}/2$, $\bm{\xi}_{1}^{\prime}=\mathbf{r}_{1}-\bm{\zeta}_{1}/2$, and $\bm{\xi}_{2}=\mathbf{r}_{2}+\bm{\zeta}_{2}/2$, $\bm{\xi}_{2}^{\prime}=\mathbf{r}_{2}-\bm{\zeta}_{2}/2$, with $\mathrm{d}^{3}\xi_{1}\,\mathrm{d}^{3}\xi_{2}\,\mathrm{d}^{3}\xi_{1}^{\prime}\,\mathrm{d}^{3}\xi_{2}^{\prime}=\,\mathrm{d}^{3}r_{1}\,\mathrm{d}^{3}r_{2}\,\mathrm{d}^{3}\zeta_{1}\,\mathrm{d}^{3}\zeta_{2}$.
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{9}(2\pi\sigma^{2})^{9/2}}\,\frac{8\sqrt{2}\pi^{3/2}{|J^{-1}|^{3}}\sigma^{3}(2+\kappa)^{3}}{\left(2+\kappa^{2}\right)^{3/2}}\,\int\,\mathrm{d}^{3}r_{1}\,\mathrm{d}^{3}r_{2}\,\mathrm{d}^{3}\zeta_{1}\,\mathrm{d}^{3}\zeta_{2}\,\,\mathrm{d}^{3}q_{1}\,\mathrm{d}^{3}q_{2}\,\,\\ \times\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\mathbf{r}_{1}+\frac{\bm{\zeta}_{1}}{2},\mathbf{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\mathbf{r}_{1}-\frac{\bm{\zeta}_{1}}{2},\mathbf{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\mathbf{q}_{1}\cdot\bm{\zeta}_{1}+\mathbf{q}_{2}\mathbf{\cdot}\bm{\zeta}_{2})}e^{-\frac{r_{1}^{2}(2\kappa+1)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}$ (88)
@@
$\mathcal{D}(\mathbf{q}_{1},\mathbf{q}_{2},\mathbf{r}_{1},\mathbf{r}_{2})=\int\mathrm{d}^{3}\zeta_{1}\int\mathrm{d}^{3}\zeta_{2}\,\,\,\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\mathbf{r}_{1}+\frac{\bm{\zeta}_{1}}{2},\mathbf{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\mathbf{r}_{1}-\frac{\bm{\zeta}_{1}}{2},\mathbf{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\mathbf{q}_{1}\cdot\bm{\zeta}_{1}+\mathbf{q}_{2}\mathbf{\cdot}\bm{\zeta}_{2})}$ (89)
Within the approach developed by Congleton et al [41], the wave function $\varphi_{{{}^{3}_{\Lambda}{\rm H}}}$ is defined as the product of $\Lambda$–d wave function and deuteron wave function, assuming deuteron is undisturbed bound object at the core.
@@
$\mathcal{D}(\mathbf{q}_{1},\mathbf{q}_{2},\mathbf{r}_{1},\mathbf{r}_{2})$ $=\int\mathrm{d}^{3}\zeta_{1}\int\mathrm{d}^{3}\zeta_{2}\,\,\,\varphi_{\Lambda}(\mathbf{r}_{1}+\frac{\bm{\zeta}_{1}}{2})\varphi^{*}_{\Lambda}(\mathbf{r}_{1}-\frac{\bm{\zeta}_{1}}{2}),\varphi_{d}(\mathbf{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\varphi^{*}_{d}(\mathbf{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\mathbf{q}_{1}\cdot\bm{\zeta}_{1}+\mathbf{q}_{2}\mathbf{\cdot}\bm{\zeta}_{2})}$ (91)
$=\int\mathrm{d}^{3}\zeta_{1}\,\,\varphi_{\Lambda}(\mathbf{r}_{1}+\frac{\bm{\zeta}_{1}}{2})\varphi^{*}_{\Lambda}(\mathbf{r}_{1}-\frac{\bm{\zeta}_{1}}{2})e^{i\mathbf{q}_{1}\cdot\bm{\zeta}_{1}}\,\int\mathrm{d}^{3}\bm{\zeta}_{2}\,\varphi_{d}(\mathbf{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\varphi^{*}_{d}(\mathbf{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i\mathbf{q}_{2}\mathbf{\cdot}\bm{\zeta}_{2}}$
$=\mathcal{D}_{\Lambda}(\mathbf{q}_{1},\mathbf{r}_{1})\mathcal{D}_{d}(\mathbf{q}_{2},\mathbf{r}_{2})$
Where, the Wigner densities $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$, is computed in coordinate space. The computation of $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$ has already been provided in a previous work [27], using the Argonne $\text{Argonne }v_{18}$ wave function. While $\mathcal{D}_{\Lambda}(\mathbf{q}_{1},\mathbf{r}_{1})$ can be easily computed using the momentum-space wave function by the Congleton approach
@@
$N(Q)=\left\{\frac{\pi}{4\alpha}\left[\left(\frac{4\alpha^{2}}{Q^{2}}+1\right)\text{Cerfe}\left(\frac{\sqrt{2}\alpha}{Q}\right)-\frac{2\alpha\left(\frac{2}{\pi}\right)^{1/2}}{Q}\right]\right\}^{-1/2}$ (93)
For $\alpha=0.068$ fm-1 $N^{2}(2.5\leavevmode\nobreak\ \mathrm{fm}^{-1})=0.094$ and $\int|\phi_{\Lambda}(q)|^{2}\mathrm{d}^{3}q=4\pi$. Finally, the Wigner density of the hypertriton $\Lambda$ part is given by
[…]
and $\mathcal{D}_{\Lambda}(0,0)/4\pi=8.0$ as expected. The probability is given as
$\mathcal{P}(q_{1},q_{2},\sigma)$ $=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}|J|^{-3}\,\frac{(2+\kappa)^{3}}{(2+\kappa^{2})^{3/2}}\,\int\mathrm{d}^{3}r_{1}\int\mathrm{d}^{3}r_{2}\,\,\mathcal{D}_{\Lambda}(\mathbf{q}_{1},\mathbf{r}_{1})\mathcal{D}_{d}(\mathbf{q}_{2},\mathbf{r}_{2})e^{-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}$ (95)
$=\frac{\left(2+\kappa^{2}\right)^{3/2}}{(2+\kappa)^{3}}\times\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}\,\,\int\mathrm{d}^{3}r_{1}\int\mathrm{d}^{3}r_{2}\,\,\mathcal{D}_{\Lambda}(\mathbf{q}_{1},\mathbf{r}_{1})\mathcal{D}_{d}(\mathbf{q}_{2},\mathbf{r}_{2})e^{-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}$
with $\int\mathrm{d}^{3}q_{1}\,\mathrm{d}^{3}q_{2}\,\frac{|J|^{3}}{(2\pi)^{9}}\,G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\mathbf{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}=1$.

### v2

@@ Appendix G Congleton Formalism for Hypertriton
$\bm{p}_{1}$ $=\frac{\kappa(2+\kappa)\bm{P}}{2+\kappa^{2}}+\frac{(2+\kappa)\bm{q}_{2}}{2+\kappa^{2}}\,$ (66)
$\bm{p}_{2}$ $=\frac{(2+\kappa)\bm{P}}{2+\kappa^{2}}+\bm{q}_{1}-\frac{\kappa(2+\kappa)\bm{q}_{2}}{2(2+\kappa^{2})},$
$\bm{p}_{3}$ $=\frac{(2+\kappa)\bm{P}}{2+\kappa^{2}}-\bm{q}_{1}-\frac{\kappa(2+\kappa)\bm{q}_{2}}{2(2+\kappa^{2})}\ .$
@@
With the Jacobian of the transformations, $|J|=-\frac{(2+\kappa)^{2}}{2+\kappa^{2}}$, the coordinates $\bm{x}_{i},\bm{p}_{i}$ satisfies the following condition: $\bm{p}_{1}\cdot\bm{x}_{1}+\bm{p}_{2}\cdot\bm{x}_{2}+\bm{p}_{3}\cdot\bm{x}_{3}=\bm{\xi}_{1}\cdot\bm{q}_{1}+\bm{\xi}_{2}\cdot\bm{q}_{2}+\bm{R}\cdot\bm{P}$. Finally defining $\bm{R}=\frac{\bm{\zeta}_{1}+\bm{\zeta}_{2}}{2}$ and $\bm{R}^{\prime}=\frac{\bm{\zeta}_{2}-\bm{\zeta}_{1}}{2}$ as well as $\mathrm{d}^{3}\bm{x}_{1}\,\mathrm{d}^{3}\bm{x}_{2}\,\mathrm{d}^{3}\bm{x}_{3}\,\mathrm{d}^{3}\bm{x}_{1}^{\prime}\,\mathrm{d}^{3}\bm{x}_{2}^{\prime}\,\mathrm{d}^{3}\bm{x}_{3}^{\prime}=\frac{|J^{-1}|^{6}}{8}\,\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\,\mathrm{d}^{3}\bm{\xi}_{1}^{\prime}\,\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\,\mathrm{d}^{3}\bm{\zeta}_{1}\,\mathrm{d}^{3}\bm{\zeta}_{2}$ with $\mathrm{d}^{3}\bm{R}\,\mathrm{d}^{3}\bm{R}^{\prime}=\frac{1}{8}\,\mathrm{d}^{3}\bm{\zeta}_{1}\,\mathrm{d}^{3}\bm{\zeta}_{2}$, we obtain the yield in terms of the Jacobi coordinates and the Hhpertriton wave function as
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{12}}\left(\frac{|J^{-1}|}{2}\right)^{3}\,\int\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\,\mathrm{d}^{3}\bm{\xi}_{1}^{\prime}\,\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\,\mathrm{d}^{3}\bm{\zeta}_{1}\,\mathrm{d}^{3}\bm{\zeta}_{2}\,\,\mathrm{d}^{3}\bm{q}_{1}\,\mathrm{d}^{3}\bm{q}_{2}\,\mathrm{d}^{3}\bm{P}\,\,e^{i\bm{\zeta}_{1}\cdot(\bm{P}-\bm{P}_{{{}^{3}_{\Lambda}{\rm H}}})}e^{-\frac{\left(2+\kappa^{2}\right)\zeta_{2}^{2}}{8\sigma^{2}(2+\kappa)^{2}}}\\ \times\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\,\,e^{i(\bm{q}_{1}\cdot\bm{\xi}_{1}-\bm{q}_{1}\cdot\bm{\xi}_{1}^{\prime}+\bm{q}_{2}\bm{\cdot}\bm{\xi}_{2}-\bm{q}_{2}\cdot\bm{\xi}_{2}^{\prime})}\times(2\pi\sigma^{2})^{-9/2}e^{\chi^{2}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}$ (67)
with a substitution of $\chi^{2}$ defined as
$\chi^{2}=-\frac{4\left((\xi_{1}+\xi_{1}^{\prime})^{2}+2(\xi_{2}+\xi_{2}^{\prime})^{2}\right)+4\kappa(\xi_{1}+\xi_{1}^{\prime})^{2}+\kappa^{2}\left(\xi_{1}^{2}+\bm{\xi}_{1}^{\prime}\cdot(2\bm{\xi}_{1}+\bm{\xi}_{1}^{\prime})+4\xi_{2}^{2}+4\bm{\xi}_{2}^{\prime}\cdot(2\bm{\xi}_{2}+\bm{\xi}_{2}^{\prime})\right)}{16\sigma^{2}(2+\kappa)^{2}}\ .$ (68)
@@
Performing pure $\delta$ integral $\int\mathrm{d}^{3}\zeta_{1}\,e^{i(\bm{P}-\bm{P}_{{{}^{3}_{\Lambda}{\rm H}}})\cdot\bm{\zeta}_{1}}=(2\pi)^{3}\delta(\bm{P}-\bm{P}_{{{}^{3}_{\Lambda}{\rm H}}})$ and $\int\mathrm{d}^{3}P$ and $\int\mathrm{d}^{3}\zeta_{2}$ in Eq. (67), we obtain
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{9}(2\pi\sigma^{2})^{9/2}}\,\frac{8\sqrt{2}\pi^{3/2}{|J^{-1}|^{3}}\sigma^{3}(2+\kappa)^{3}}{\left(2+\kappa^{2}\right)^{3/2}}\,\int\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\,\mathrm{d}^{3}\bm{\xi}_{1}^{\prime}\,\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}\,\,\mathrm{d}^{3}\bm{q}_{1}\,\mathrm{d}^{3}\bm{q}_{2}\,\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1},\bm{\xi}_{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{\xi}_{1}^{\prime},\bm{\xi}_{2}^{\prime})\\ \times e^{i(\bm{q}_{1}\cdot(\bm{\xi}_{1}-\bm{\xi}_{1}^{\prime})+\bm{q}_{2}\cdot(\bm{\xi}_{2}-\bm{\xi}_{2}^{\prime}))}\times e^{\chi^{2}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}$ (69)
@@
Further simplifications are made by making the following coordinate transformations $\bm{\xi}_{1}=\bm{r}_{1}+\bm{\zeta}_{1}/2$, $\bm{\xi}_{1}^{\prime}=\bm{r}_{1}-\bm{\zeta}_{1}/2$, and $\bm{\xi}_{2}=\bm{r}_{2}+\bm{\zeta}_{2}/2$, $\bm{\xi}_{2}^{\prime}=\bm{r}_{2}-\bm{\zeta}_{2}/2$, with $\mathrm{d}^{3}\bm{\xi}_{1}\,\mathrm{d}^{3}\bm{\xi}_{2}\,\mathrm{d}^{3}\bm{\xi}_{1}^{\prime}\,\mathrm{d}^{3}\bm{\xi}_{2}^{\prime}=\,\mathrm{d}^{3}\bm{r}_{1}\,\mathrm{d}^{3}\bm{r}_{2}\,\mathrm{d}^{3}\bm{\zeta}_{1}\,\mathrm{d}^{3}\bm{\zeta}_{2}$. In this way we obtain
@@
$\frac{\mathrm{d}N_{{{}^{3}_{\Lambda}{\rm H}}}}{\mathrm{d}^{3}P}=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{9}(2\pi\sigma^{2})^{9/2}}\,\frac{8\sqrt{2}\pi^{3/2}{|J^{-1}|^{3}}\sigma^{3}(2+\kappa)^{3}}{\left(2+\kappa^{2}\right)^{3/2}}\,\int\,\mathrm{d}^{3}\bm{r}_{1}\,\mathrm{d}^{3}\bm{r}_{2}\,\mathrm{d}^{3}\bm{\zeta}_{1}\,\mathrm{d}^{3}\bm{\zeta}_{2}\,\,\mathrm{d}^{3}\bm{q}_{1}\,\mathrm{d}^{3}\bm{q}_{2}\,\,\\ \times\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{r}_{1}+\frac{\bm{\zeta}_{1}}{2},\bm{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{r}_{1}-\frac{\bm{\zeta}_{1}}{2},\bm{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\bm{q}_{1}\cdot\bm{\zeta}_{1}+\bm{q}_{2}\bm{\cdot}\bm{\zeta}_{2})}e^{-\frac{r_{1}^{2}(\kappa+2)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}\ ,$ (70)
@@
$\mathcal{D}(\bm{q}_{1},\bm{q}_{2},\bm{r}_{1},\bm{r}_{2})=\int\mathrm{d}^{3}\bm{\zeta}_{1}\int\mathrm{d}^{3}\bm{\zeta}_{2}\,\,\,\varphi_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{r}_{1}+\frac{\bm{\zeta}_{1}}{2},\bm{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\,\varphi^{*}_{{{}^{3}_{\Lambda}{\rm H}}}(\bm{r}_{1}-\frac{\bm{\zeta}_{1}}{2},\bm{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\bm{q}_{1}\cdot\bm{\zeta}_{1}+\bm{q}_{2}\bm{\cdot}\bm{\zeta}_{2})}\ .$ (71)
Within the approach developed in Ref. [34], the wave function $\varphi_{{{}^{3}_{\Lambda}{\rm H}}}$ is defined as the product of $\Lambda$–d wave function and deuteron wave function, assuming deuteron is undisturbed bound object at the core. Therefore
@@
$\mathcal{D}(\bm{q}_{1},\bm{q}_{2},\bm{r}_{1},\bm{r}_{2})$ $=\int\mathrm{d}^{3}\bm{\zeta}_{1}\int\mathrm{d}^{3}\bm{\zeta}_{2}\,\,\,\varphi_{\Lambda}(\bm{r}_{1}+\frac{\bm{\zeta}_{1}}{2})\varphi^{*}_{\Lambda}(\bm{r}_{1}-\frac{\bm{\zeta}_{1}}{2}),\varphi_{d}(\bm{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\varphi^{*}_{d}(\bm{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i(\bm{q}_{1}\cdot\bm{\zeta}_{1}+\bm{q}_{2}\bm{\cdot}\bm{\zeta}_{2})}$ (73)
$=\int\mathrm{d}^{3}\bm{\zeta_{1}}\,\,\varphi_{\Lambda}(\bm{r}_{1}+\frac{\bm{\zeta}_{1}}{2})\varphi^{*}_{\Lambda}(\bm{r}_{1}-\frac{\bm{\zeta}_{1}}{2})e^{i\bm{q}_{1}\cdot\bm{\zeta}_{1}}\,\int\mathrm{d}^{3}\bm{\zeta}_{2}\,\varphi_{d}(\bm{r}_{2}+\frac{\bm{\zeta}_{2}}{2})\varphi^{*}_{d}(\bm{r}_{2}-\frac{\bm{\zeta}_{2}}{2})e^{i\bm{q}_{2}\bm{\cdot}\bm{\zeta}_{2}}$
$=\mathcal{D}_{\Lambda}(\bm{q}_{1},\bm{r}_{1})\mathcal{D}_{d}(\bm{q}_{2},\bm{r}_{2})\ ,$
where the Wigner densities $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$ is computed in coordinate space. The computation of $\mathcal{D}_{d}(\bm{k}_{2},\bm{r}_{2})$ has already been provided in a previous work [41], using the $\text{Argonne }v_{18}$ wave function, while $\mathcal{D}_{\Lambda}(\bm{q}_{1},\bm{r}_{1})$ can be easily computed using the momentum-space wave function by the Congleton approach
@@
$N(Q)=\left\{\frac{\pi}{4\alpha}\left[\left(\frac{4\alpha^{2}}{Q^{2}}+1\right)\text{Cerfe}\left(\frac{\sqrt{2}\alpha}{Q}\right)-\frac{2\alpha\left(\frac{2}{\pi}\right)^{1/2}}{Q}\right]\right\}^{-1/2}\ .$ (75)
Here $\alpha=0.068$ fm-1, $N^{2}(2.5\penalty 10000\ \mathrm{fm}^{-1})=0.094$ and $\int|\phi_{\Lambda}(q)|^{2}\mathrm{d}^{3}\bm{q}=4\pi$. Finally, the Wigner density of the hypertriton $\Lambda$ part is given by
$\mathcal{D}_{\Lambda}(\bm{q},\bm{r})=N(Q)^{2}\int_{0}^{\infty}dk\frac{4i\pi\left(-1+e^{2ikr}\right)}{qr\left(k^{2}+4\left(\alpha^{2}+q^{2}\right)\right)}\log\left(\frac{4\alpha^{2}+(k-2q)^{2}}{4\alpha^{2}+(k+2q)^{2}}\right)e^{-\frac{k^{2}+2ikQ^{2}r+4q^{2}}{2Q^{2}}}\ ,$ (76)
and $\mathcal{D}_{\Lambda}(0,0)/4\pi=8.0$ as expected. The probability is finally written as
$\mathcal{P}(q_{1},q_{2},\sigma)$ $=\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}|J|^{-3}\,\frac{(2+\kappa)^{3}}{(2+\kappa^{2})^{3/2}}\,\int\mathrm{d}^{3}\bm{r}_{1}\int\mathrm{d}^{3}\bm{r}_{2}\,\,\mathcal{D}_{\Lambda}(\bm{q}_{1},\bm{r}_{1})\mathcal{D}_{d}(\bm{q}_{2},\bm{r}_{2})e^{-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}$ (77)
$=\frac{\left(2+\kappa^{2}\right)^{3/2}}{(2+\kappa)^{3}}\times\frac{S_{\mathrm{{{}^{3}_{\Lambda}{\rm H}}}}}{(2\pi)^{3}\sigma^{6}}\,\,\int\mathrm{d}^{3}\bm{r}_{1}\int\mathrm{d}^{3}\bm{r}_{2}\,\,\mathcal{D}_{\Lambda}(\bm{q}_{1},\bm{r}_{1})\mathcal{D}_{d}(\bm{q}_{2},\bm{r}_{2})e^{-\frac{r_{1}^{2}(2+\kappa)^{2}+4r_{2}^{2}\left(2+\kappa^{2}\right)}{4\sigma^{2}(2+\kappa)^{2}}}$
with $\int\mathrm{d}^{3}\bm{q}_{1}\,\mathrm{d}^{3}\bm{q}_{2}\,\frac{|J|^{3}}{(2\pi)^{9}}\,G_{\mathrm{np}\Lambda}(\bm{p}_{1},\bm{p}_{2},\bm{p}_{3})|_{{\bm{P}}={\bm{P}}_{{}^{3}_{\Lambda}\text{H}}}=1$.